OpenStudy (anonymous):

Let f(x) = (x^2+3x-4) and g(x) = (x+4) Find f*g and state the domain. Find f/g and state the domain. Help! I need help solving, I think I know the first one just need some advice, but the second one I have no clue!!!!!

OpenStudy (anonymous):

First one, (f*g)(x) = fx*gx= (x^2+3x-4)(x+4) x(x^2+3x-4)+4(x^2+3x-4) x^3+3x^2-4x+4x^2+12x-16 x^3+7x^2+8x-16

OpenStudy (ranga):

For f*g , just multiply the two functions: (x^2+3x-4)(x+4) The first expression can be factored: (x^2+3x-4) = (x+4)(x-1) f*g = (x+4)(x-1)(x+4) = (x-1)(x+4)^2 You can leave it as factors or multiply them out. Domain means what are the allowed values of x. Here we see no restrictions on x. x can be any real number. So the domain is -infinity < x < infinity or in interval notation: (-infinity, infinity).

OpenStudy (ranga):

f/g = (x^2+3x-4) / (x+4) = (x+4)(x-1) / (x+4) Since we cannot divide by zero, x cannot be -4. So the domain must exclude -4 Domain is: (-infinity, -4) union (-4, infinity) You may be tempted to cancel out the (x+4) in the numerator and denominator and say there is no restriction on the domain but that would be incorrect. When x = -4 it is dividing 0/0 and you can't cancel out the zeros. But for all x not equal to -4 f/g = (x-1) (very important to remember that x cannot be -4).

OpenStudy (anonymous):

So is my first answer in the ballpark?

OpenStudy (ranga):

Your first answer is correct. Just include the domain.

OpenStudy (anonymous):

K, how would approach solving the second?

OpenStudy (anonymous):

(f/g)(x) = (x^2+3x-4)/x+4

OpenStudy (ranga):

Yes and the domain should exclude -4. See my answer above where I have solved the second question.

OpenStudy (anonymous):

So it can't be reduced further

OpenStudy (ranga):

You can say either way: f/g = (x-1) where x is not equal to -4 or the domain is (-infinity, -4) U (-4, infinity) or f/g = (x^2+3x-4)/(x+4) where x is not equal to -4 or the domain is (-infinity, -4) U (-4, infinity) I do not know which form your teacher or text book prefers. I will look for similar examples in the book and express it the same way. But both are correct. The first one is nicely simplified.

OpenStudy (anonymous):

so second answer would be (x+4)(x-1) / (x+4) and the domain is all real number except x (not equal to) -4

OpenStudy (ranga):

Yes. But if your book prefers the simplified form then the second answer will be: (x-1) for all real number for x except -4. If the function f/g were plotted it will look like this (a gap or hole or discontinuity at x = -4):

OpenStudy (anonymous):

yes simplified version is the best, so I get the domain part but what is f/g ?

OpenStudy (ranga):



OpenStudy (anonymous):

Right so the first one simplifies to x^3+7x^2+8x-16 what does the second one simplify too?

OpenStudy (ranga):

We have been discussing the second solution for quite a while now. The last few replies are all for the second question.

OpenStudy (anonymous):

Right I understand that, especially the domain aspect I just wasn't sure other than factoring it to (x+4)(x-1) / (x+4), if that's all i needed to do to find f/g

OpenStudy (ranga):

After factoring you cancel out the (x+4) at the top and the bottom. That leaves with just (x-1) at the top. f/g(x) = x -1 for all x not equal to -4.

OpenStudy (anonymous):

Got it thanks! sorry i little slow, my teacher often gets impatient haha

OpenStudy (ranga):

np. You are welcome.